An IV is infusing at 35 gtt/min. The drop factor is 15 gtt/mL. In 5 hours, the patient will have received how many mLs?
Answer #1

ANSWER.

Calculation of volume(ml) of drug required.

       gtt/min = Volume(ml) /Time(min) * Drop factor(gtt/ml)

So

   Formula

   Volume (ml) = (gtt/min) / drop factor * Time(min)

Here

        gtt/min = 35 gtt/min.

        Drop factor = 15 gtt/ml

         Time = 300 min (5hr)

As per above formula for volume

               Volume(ml) = 35 /15 * 300

                                   = 2.33 * 300

                                   = 699 ml